A body starts moving with velocity 10 m/s. It ex experience a retardation equal to 0.2v^2. Its velocity after 2s is

Grade:11

1 Answers

Susmita
5 years ago
Retardation
f=-0.2v2
or,dv/dt=-0.2v2
or,dv/v2=-0.2dt
Integrate for dv from 10 to v and dt from 0 to 2.
(1/v)-(1/10)=0.2×2=0.4
Or,1/v=0.4+0.1=0.5
Or,v=2 m/s
Hope this helps.